OpenStudy (anonymous):

Find the solutions to this equation algebraically using the Fundamental Theorem of Algebra, the Rational Root Theorem, Descartes' Rule of Signs, and the Factor Theorem: V = x^3 - x^2 - 20x

OpenStudy (anonymous):

Factor Theorem \[V = x^3 - x^2 - 20x=x(x^2-x-20)=x(x^2-5x+4x-20)\\=x(x(x-5)+4(x-5))=x(x-5)(x+4)\]

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

welcome